OpenStudy (anonymous):

What degree Fahrenheit is it if the temperature is 15 degrees Celsius? Show all steps or reasoning

OpenStudy (anonymous):

I assume we can use the formula: F = 9c/5 + 32 c = 15, so plug in the equation. F = 9 (15)/5 + 32 F = 27 + 32 F 59